What is the value of x in the solutions to following equations
x - y + -3
x + 3y =5

Answer by ewatrrr(24785)

 
Hi
Solution to the following system:
x - y = -3 **(x = y-3)
x + 3y = 5
-x + y = 3 ||||multiplying 1st EQ thru by -1 to eliminate the x variable
x + 3y = 5
4y = 8
y = 2 and x = -1 (x = y-3)
Ordered pair (-1,2) is the solution for this system